# Impact of Small-Scale Gravity Waves on Tracer Transport

### v1

@@ 3 Summary of equations in dimensional form
    Solving for the leading-order buoyancy amplitude $b_{\alpha 1}^{\prime}=\left|b_{\alpha 1}^{\prime}\right|e^{i\varLambda_{\alpha}}$ from the wave energy equation
@@
    Integrating the prognostic equation for $\varLambda_{\alpha}$ to obtain its large-scale phase:
    $\begin{split}\left(\partial_{t}+\mathbf{c}_{g,\alpha}\cdot\nabla\right)\varLambda_{\alpha}=&-\frac{fm_{\alpha}}{2\left|\mathbf{k}_{\alpha}\right|^{2}\hat{\omega}_{\alpha}}\left(\mathbf{k}_{\alpha}\times\nabla\right)\left\langle{\mathbf{u}}\right\rangle\\ &+\frac{N^{2}}{2\left(N^{2}-f^{2}\right)m_{\alpha}}\left[c_{p}\overline{\theta}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla\right)\left\langle{\pi}\right\rangle+\frac{1}{\overline{\theta}N^{2}}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla\right)\left\langle{\theta}\right\rangle\right]\\ &-\frac{fm_{\alpha}^{2}}{\left|\mathbf{k}_{\alpha}\right|^{4}\hat{\omega}_{\alpha}}\left[\left(k_{\alpha}\partial_{y}+l_{\alpha}\partial_{x}\right)\hat{\omega}_{\alpha}+l_{\alpha}\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)k_{\alpha}-k_{\alpha}\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)l_{\alpha}\right]\;.\end{split}$ (68)
@@
    Using the polarization relations (Equations 54, 55 and 56) to compute the leading-order wind and pressure amplitudes.
@@
    Determining the right-hand sides (Equations 60, 61, 62 and 63 of Equation 58) of Equation 58 based on the leading-order wave amplitudes.
@@
    Solving Equation 58 for the next-order wind amplitudes.
@@
    Computing the leading-order tracer amplitude from Equation 57 and the next-order tracer amplitude from Equation 64.
@@
    Finally, substituting these values, along with the leading-order and next-order wind amplitudes, into Equation 66 to obtain $\mathcal{Q}^{(1)}$.
Since $\det(M_{\alpha})=0$ (leading to the dispersion relation $\hat{\omega}_{\alpha}$ in Equation 52), the matrix $M_{\alpha}$ is not invertible. However, by computing the pseudo-inverse $M_{\alpha}^{-1}$, as described by Golub and Kahan [11], we obtain an approximate solution for $\mathbf{Z}_{\alpha}^{\prime\prime}$ as
$\mathbf{Z}_{\alpha}^{\prime\prime}\approx M_{\alpha}^{-1}\mathbf{R}_{\alpha}\;.$ (69)
@@ 4 Numerical Simulation of Gravity Wave-Induced Tracer Transport
We extend a gravity wave parameterization model to compute the previously discussed gravity wave tracer flux convergences. The model is coupled with a flow solver, both of which are introduced in Section 4.1. The treatment of leading-order and next-order flux convergences is detailed in Sections 4.2 and 4.3, respectively. To validate our approach, we compare parameterized gravity wave simulations at coarse resolution against high-resolution, wave-resolving reference simulations.
@@ 4.1 Numerical Model
In the following section, we introduce the Pseudo-Incompressible Flow solver (PincFlow) used to simulate the resolved flow. PincFlow operates in two modes: a high-resolution, wave-resolving mode for generating reference data and a low-resolution mode with parameterized gravity waves, incorporating a Lagrangian ray tracer. We then present the gravity wave model, which computes the gravity wave impact on the resolved flow and is couple to PincFlow.
@@ 4.1.1 PincFlow
PincFlow solves the pseudo-incompressible equations, first introduced by [9], in conservative flux form, as presented by Klein [13], Rieper et al. [16], Schmid et al. [17], and Jochum et al. [12]. The model includes a passive tracer mixing ratio, $\psi$, whose transport is described by Equation 1 and, in flux form, by
@@
Atmospheric variables are discretized on a C-grid [4], where density, Exner pressure, and tracer mixing ratio are defined at cell centers, while the wind components $u$, $v$, and $w$ are staggered along the $x$-, $y$-, and $z$-directions, respectively. Mass, tracer, and momentum fluxes are computed using a monotone upwind scheme [18]. The equations are integrated using a semi-implicit time-stepping scheme [17]. The tracer advection equation (Equation 70) is solved explicitly twice per time step, each over half a step. For explicit time integration, the third-order Runge-Kutta method of Williamson [22] is employed.
@@ 4.1.2 PincFlow/MS-GWaM
In the next-two sections, we extend MS-GWaM to calculate the leading- and next-order gravity wave tracer flux convergences, as discussed in Sections 2 and 3. To validate these extensions, we use two test cases of gravity wave packets, each designed so that one of the flux convergences dominates the gravity wave forcing on the large-scale tracer.
@@ 4.2 Parameterization of the Leading-order Gravity Wave Tracer Flux Convergence
In this section, we present the parameterization of the leading-order gravity wave tracer flux convergence using MS-GWaM. To validate this extension, we simulate a three-dimensional gravity wave packet propagating through a large-scale tracer field that initially varies only in the vertical direction. We then compare the resulting changes in the large-scale tracer distribution with those from a wave-resolving reference simulation.
To adapt the theory described above to the spectral Lagrangian approach of MS-GWaM, we rewrite the leading-order flux convergence in terms of $\mathcal{N}$. From Equation 65, we obtain
@@
It is important to note that the derivation of Equation 72 assumes geostrophic and hydrostatic equilibrium of the large-scale flow to ensure the numerical feasibility of simulation $\mathcal{N}$, as presented in Achatz et al. [1]. However, this assumption is not required for the derivation of $\mathcal{Q}^{(0)}$ or $\mathcal{Q}^{(1)}$. The geostrophic and hydrostatic equilibrium approximation introduces an $\mathcal{O}(\varepsilon)$ error in the prognostic equation (72) for $\mathcal{N}$. As demonstrated by Bölöni et al. [7], Wei et al. [20], and Jochum et al. [12], and as will be confirmed by the following simulation results, this error appears to be sufficiently small for the predicted $\mathcal{N}$ to yield reliable results.
The impact of gravity waves on tracer transport is incorporated by extending Equation 70 to
@@
While Equation 70 is used in wave-resolving reference simulations, the extend form, Equation 78, is applied when the model resolution is coarse and gravity waves are parameterized using MS-GWaM.
To validate the extension of PincFlow/MS-GWaM with Equations 77 and 78, we conduct three simulations of a three-dimensional gravity wave packet. A high-resolution wave-resolving simulation serves as our reference and will be referred to as 3DWR. To ensure the gravity wave is well resolved, we use at least 16 grid points per initial wavelength in each horizontal direction and 10 grid points per wavelength in the vertical direction. Additionally, we perform a coarse-resolution where the resolution is approximately one grid point per wavelength. Since the gravity wave impact, including $\mathcal{Q}^{(0)}$, is not resolved at this resolution, it is parameterized using MS-GWaM. We denote this simulation as 3DCR. Finally, to assess the effect of the tracer flux convergence in Equation 78, we conduct another simulation with parameterized waves, but with $\mathcal{Q}^{(0)}$ neglected. This simulation is referred to as 3DCRno.
The atmosphere in the simulations is initialized as follows: We begin with a rotating, isothermal atmosphere at rest, with a temperature of $T_{0}=300\text{K}$. A locally monochromatic wave packet is introduced, where the atmospheric variables are given by
@@
with wave amplitude $\hat{b}$ and initial wavenumbers $l_{0}$ and $m_{0}$ in the $y$- and $z$-directions, respectively. The initial perturbations in the wind $\mathbf{v}^{\prime}$ and tracer mixing ration $\psi^{\prime}$ are determined from $b^{\prime}$ using the polarization relations (Equations 28, 29 and 17). The buoyancy wave amplitude is initially defined by a Gaussian function in the $z$-direction and a cosine function in the $x$-direction, with no dependence on $y$:
@@
where $\alpha_{\psi}$ is an arbitrary scaling parameter. All simulation parameters are summarized in Table 2.
Table 2: Initial model parameters for the simulations of the three-dimensional wave packet (3D) and the two-dimensional wave packet (2D). The number of grid points $N_{x}$, $N_{y}$, and $N_{z}$ are given for the wave-resolving simulations (WR) and the coarse-resolution simulations with parameterized gravity waves (CR).
Parameter 3D 2D
Wavelength in $x$-direction $\lambda_{x}$ [km] $1$
Wavelength in $y$-direction $\lambda_{y}$ [km] $300$
Wavelength in $z$-direction $\lambda_{z}$ [km] $-1$ $1$
Branch of intrinsic frequency $\hat{\omega}$ $+1$ $-1$
Wave amplitude factor $a_{0}$ $0.5$ $0.1$
Width of wave packet in $x$-direction $\sigma_{x}$ [km] $1500$
Width of wave packet in $z$-direction $\sigma_{z}$ [km] $30$ $2$
Wave packet position in $x$-direction $x_{0}$ [km] $4500$
Wave packet position in $z$-direction $z_{0}$ [km] $30$ $10$
Coriolis parameter $f$ [s-1] $10^{-4}$ 00
Brunt-Vaisala frequency $N$ [s-1] $0.02$ $0.02$
Atmosphere temperature $T_{0}$ [K] 300 300
Tracer mixing ratio factor $\alpha_{\psi}$ [m-1] 1 1
Domain size in $x$-direction $L_{x}$ [km] $9000$ $1$
Domain size in $y$-direction $L_{y}$ [km] $300$
Domain size in $z$-direction $L_{z}$ [km] $100$ $30$
Number of grid points in $x$-direction (WR) $N_{x}$ $512$ 32
Number of grid points in $y$-direction (WR) $N_{y}$ $16$
Number of grid points in $z$-direction (WR) $N_{z}$ $1000$ 960
Number of grid points in $x$-direction (CR) $N_{x}$ $23$
Number of grid points in $y$-direction (CR) $N_{y}$ $1$
Number of grid points in $z$-direction (CR) $N_{z}$ $100$ 300
In the coarse-resolution simulations 3DCR and 3DCRno, the large-scale tracer distribution $\left\langle{\psi}\right\rangle$ and velocity $\left\langle{\mathbf{v}}\right\rangle$are obtained directly as output. However, in the high-resolution wave-resolving simulation 3DWR, the output consists of the full variables $\psi$ and $\mathbf{v}$, which include both the large-scale and gravity wave components. Therefore, post-processing is required to extract $\left\langle{\psi}\right\rangle$ and $\left\langle{\mathbf{v}}\right\rangle$. Since the initial wave packet setup in Equation 83 assumes no dependence on the $y$-coordinate for the large-scale variables, we can extract them by averaging the full variables over one wavelength in the $y$-direction. The gravity wave components are then determined as
@@
The results after 1500 min are visualized in Figure 1, showing the gravity wave tracer flux convergence $\mathcal{Q}^{(0)}$ and the resulting change in the tracer distribution for 3DWR, 3DCR, and 3DCRno. In the reference simulation 3DWR, the gravity wave packet leads to a decrease in the tracer mixing ratio to the left and an increase to the right of the wave packet amplitude maximum. This behaviour can be explained by the initial setup of $\mathbf{k}_{0}=(0,l_{0},m_{0})$ and $\nabla\left\langle{\psi}\right\rangle=\alpha_{\psi}\mathbf{e}_{z}$, as the wave packet has changed only slightly during the 1500-minute simulation due to its relatively small vertical group velocity. Comparing 3DCR to the reference simulation 3DWR, we find that the fluxes and the resulting changes in the large-scale tracer distribution are well reproduced. To assess the effect of setting $\mathcal{Q}^{(0)}=0$, we compare 3DCR to 3DCRno, where the latter neglects $\mathcal{Q}^{(0)}$. In 3DCRno, the large-scale tracer $\left\langle{\psi}\right\rangle$ is only advected by the wave-induced large-scale wind (not shown). The resulting $\Delta\left\langle{\psi}\right\rangle$ differs significantly from that in 3DCR, and consequently from 3DWR. Thus, we conclude that the changes in $\Delta\left\langle{\psi}\right\rangle$ observed in 3DWR and 3DCR are significantly impacted by $\mathcal{Q}^{(0)}$ and not merely a result of advection due to the large-scale wind $\left\langle{\mathbf{v}}\right\rangle$.

### v2

@@ 3 Summary of equations in dimensional form
    Solving for the leading-order buoyancy amplitude $b^{\prime}=\left|b^{\prime}\right|e^{i\varLambda}$ from the wave energy equation
    $E_{w}=\frac{\overline{\rho}}{2}\left(\frac{\left|\mathbf{v}^{\prime}\right|^{2}}{2}+\frac{\left|b^{\prime}\right|^{2}}{2N^{2}}\right)=\overline{\rho}\frac{\left|b^{\prime}\right|^{2}}{2N^{2}}\frac{N^{2}k_{h}^{2}+f^{2}m^{2}}{N^{2}k_{h}^{2}}\;.$ (68)
    Note that this relation provides $|b^{\prime}|$ once the wave energy $E_{w}$ is specified; in practice $E_{w}$ must be predicted by a gravity-wave model or another energy source.
@@
    Integrating the prognostic equation for $\varLambda$ to obtain its large-scale phase:
    $\begin{split}\left(\partial_{t}+\mathbf{c}_{g}\cdot\nabla\right)\varLambda=&-\frac{fm}{2\left|\mathbf{k}\right|^{2}\hat{\omega}}\left(\mathbf{k}\times\nabla\right)\left\langle{\mathbf{u}}\right\rangle\\ &+\frac{|\mathbf{k}|^{2}}{2m\left(N^{2}-f^{2}\right)}\left[\frac{g}{\overline{\theta}}\left(\hat{\mathbf{c}}_{g}\cdot\nabla\right)\left\langle{\theta}\right\rangle-\frac{N^{2}}{g}c_{p}\overline{\theta}\left(\hat{\mathbf{c}}_{g}\cdot\nabla\right)\left\langle{\pi}\right\rangle\right]\\ &-\frac{f\hat{\omega}m^{2}}{k_{h}^{2}\left(N^{2}k_{h}^{2}+f^{2}m^{2}\right)}\left[\left(l\partial_{x}+k\partial_{y}\right)\hat{\omega}+l\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)k-k\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)l\right]\;.\end{split}$ (69)
@@
    Using the polarization relations (Equations˜55, 56 and 57) to compute the leading-order wind and pressure amplitudes.
@@
    Determining the right-hand sides (Equations˜61, 62, 63 and 64 of Equation˜59) of Equation˜59 based on the leading-order wave amplitudes.
@@
    Solving Equation˜59 for the next-order wind amplitudes.
@@
    Computing the leading-order tracer amplitude from Equation˜58 and the next-order tracer amplitude from Equation˜65.
@@
    Finally, substituting these values, along with the leading-order and next-order wind amplitudes, into Equation˜67 to obtain $\mathcal{Q}^{(1)}$.
A difficulty arises in step 5: As the dispersion relation $\hat{\omega}$ (Equation˜53) is obtained by enforcing $\det(M)=0$, the inverse of $M$ cannot be computed directly. However, it is important to note that the next-order amplitude vector $\mathbf{Z}^{\prime\prime}$ does not belong to the null space of $M$, as the latter is assumed to only contribute to the leading-order wave amplitudes $\mathbf{Z}^{\prime}$ (Equation˜32). This allows us to determine the exact solution for $\mathbf{Z}^{\prime\prime}$ by determining the inverse using the singular value decomposition [SVD; 12]. Writing
$M=U\Sigma V^{*}\;,$ (70)
with unitary matrices $U$ and $V$, and diagonal $\Sigma$ containing the singular values, the pseudo-inverse is
$M^{-1}=V\Sigma^{-1}U^{*}\;,$ (71)
where $\Sigma^{-1}$ is obtained by inverting non-zero singular values. This yields the exact solution
$\mathbf{Z}^{\prime\prime}=M^{-1}\mathbf{R}\;.$ (72)
@@ 4 Numerical Simulation of Gravity Wave-Induced Tracer Transport
We extend a gravity wave parameterization model to compute the previously discussed gravity wave tracer flux convergences. The model is coupled with a flow solver, both of which are introduced in Section˜4.1. The treatment of leading-order and next-order flux convergences is detailed in Sections˜4.2 and 4.3, respectively. To validate our approach, we compare parameterized gravity wave simulations at coarse resolution against high-resolution, wave-resolving reference simulations.
@@ 4.1 Numerical Model
In the following section, we introduce the Pseudo-Incompressible Flow solver (PincFlow) used to simulate the resolved flow. PincFlow operates in two modes: a high-resolution, wave-resolving mode for generating reference data and a low-resolution mode with parameterized gravity waves, incorporating a Lagrangian ray tracer. We then present the gravity wave model, which computes the gravity wave impact on the resolved flow and is coupled to PincFlow.
@@ 4.1.1 PincFlow
PincFlow solves the pseudo-incompressible equations, first introduced by Durran [10], in conservative flux form, as presented by Klein [15], Rieper et al. [18], Schmid et al. [19], and Jochum et al. [14]. The model includes a passive tracer mixing ratio, $\psi$, whose transport is described by Equation˜1 and, in flux form, by
@@
Atmospheric variables are discretized on a C-grid [5], where density, Exner pressure, and tracer mixing ratio are defined at cell centers, while the wind components $u$, $v$, and $w$ are staggered along the $x$-, $y$-, and $z$-directions, respectively. Mass, tracer, and momentum fluxes are computed using a monotone upwind scheme [20]. The equations are integrated using a semi-implicit time-stepping scheme [19]. The tracer advection equation (Equation˜73) is solved explicitly twice per time step, each over half a step. For explicit time integration, the third-order Runge-Kutta method of Williamson [24] is employed.
@@ 4.1.2 PincFlow/MS-GWaM
In the next-two sections, we extend MS-GWaM to calculate the leading- and next-order gravity wave tracer flux convergences, as discussed in Sections˜2 and 3. To validate these extensions, we use two test cases of gravity wave packets, each designed so that one of the flux convergences dominates the gravity wave forcing on the large-scale tracer.
@@ 4.2 Parameterization of the Leading-order Gravity Wave Tracer Flux Convergence
Figure 1: Gravity wave-induced tracer flux convergence $\mathcal{Q}^{(0)}$ (top row) and change in the large-scale tracer distribution $\Delta\left\langle{\psi}\right\rangle$ (bottom row) from a three-dimensional gravity wave packet in an atmosphere initially at rest, at $t=1500\text{min}$; (left column) 3DWR: wave-resolving reference simulation; (center column) 3DCR: coarse-resolution simulation with parameterized gravity wave and including $\mathcal{Q}^{(0)}$; (right column) 3DCRno: coarse-resolution simulation with $\mathcal{Q}^{(0)}=\mathcal{Q}^{(1)}=0$. The solid and dashed contour lines illustrate the regions with positive and negative values, respectively.
In this section, we present the parameterization of the leading-order gravity wave tracer flux convergence using MS-GWaM. To validate this extension, we conduct two simulations of three-dimensional gravity wave packets propagating through a large-scale tracer field that initially varies only in the vertical direction. We then compare the resulting changes in the large-scale tracer distribution with those from a wave-resolving reference simulation.
To adapt the theory described above to the spectral Lagrangian approach of MS-GWaM, we rewrite the leading-order flux convergence in terms of $\mathcal{N}$. From Equation˜66, we obtain
@@
It is important to note that the derivation of Equation˜75 assumes geostrophic and hydrostatic equilibrium of the large-scale flow to ensure the numerical feasibility of simulation $\mathcal{N}$, as presented in Achatz et al. [2]. However, this assumption is not required for the derivation of $\mathcal{Q}^{(0)}$ or $\mathcal{Q}^{(1)}$. The geostrophic and hydrostatic equilibrium approximation introduces an $\mathcal{O}(\varepsilon)$ error in the prognostic equation (75) for $\mathcal{N}$. As demonstrated by Bölöni et al. [8], Wei et al. [22], and Jochum et al. [14], and as will be confirmed by the following simulation results, this error appears to be sufficiently small for the predicted $\mathcal{N}$ to yield reliable results.
The impact of gravity waves on tracer transport is incorporated by extending Equation˜73 to
@@
While Equation˜73 is used in wave-resolving reference simulations, the extended form, Equation˜81, is applied when the model resolution is coarse and gravity waves are parameterized using MS-GWaM.
To validate the extension of PincFlow/MS-GWaM with Equations˜80 and 81, we conduct two test cases examining the impact of a three-dimensional gravity wave packet on the large-scale tracer distribution. In the first case (Section˜4.2.1), a gravity wave propagates through an initially resting atmosphere and has an initial aspect ratio of $\lambda_{y}/\lambda_{z}=$300$\$, where $\lambda_{y}$ and $\lambda_{z}$ are the horizontal and vertical wavelengths, respectively. The second case (Section˜4.2.2) introduces a background jet and uses a gravity wave with an initial aspect ratio $\lambda_{y}/\lambda_{z}=$10$\$. These contrasting setups illustrate that the parameterization performs consistently across different background flows and wave configurations.
In both test cases, a high-resolution wave-resolving simulation serves as our reference and will be referred to as 3DWR. To ensure the gravity wave is well resolved, we use at least 16 grid points per initial wavelength in each horizontal direction and 10 grid points per wavelength in the vertical direction. Additionally, we perform a coarse-resolution where the resolution is approximately one grid point per wavelength. Since the gravity wave impact, including $\mathcal{Q}^{(0)}$, is not resolved at this resolution, it is parameterized using MS-GWaM. We denote this simulation as 3DCR. Finally, to assess the effect of the tracer flux convergence in Equation˜81, we conduct another simulation with parameterized waves, but with $\mathcal{Q}^{(0)}$ neglected. This simulation is referred to as 3DCRno. As $\mathcal{Q}^{(0)}$ is the largest gravity wave-effect on tracer transport in these cases, both 3DCR and 3DCRno do not include the next-order effect $\mathcal{Q}^{(1)}$.
4.2.1 Wave packet in an atmosphere at rest
The first case we present is of a gravity wave packet in a rotating, isothermal atmosphere at rest, with a temperature of $T_{0}=300\text{K}$. A locally monochromatic wave packet is introduced, where the atmospheric variables are given by
@@ 4.2.1 Wave packet in an atmosphere at rest
with wave amplitude $\hat{b}$ and initial wavenumbers $l_{0}$ and $m_{0}$ in the $y$- and $z$-directions, respectively. The initial perturbations in the wind $\mathbf{v}^{\prime}$ and tracer mixing ration $\psi^{\prime}$ are determined from $b^{\prime}$ using the polarization relations (Equations˜29, 30 and 18). The buoyancy wave amplitude is initially defined by a Gaussian function in the $z$-direction and a cosine function in the $x$-direction, with no dependence on $y$:
@@
where $\alpha_{\psi}$ is an arbitrary scaling parameter. All simulation parameters are summarized in Table˜2.
Table 2: Initial model parameters for the simulations of the three-dimensional wave packet in an initially resting atmosphere (3D), in an atmosphere with a jet (3Djet), and the two-dimensional wave packet (2D). The number of grid points $N_{x}$, $N_{y}$, and $N_{z}$ are given for the wave-resolving simulations (WR) and the coarse-resolution simulations with parameterized gravity waves (CR).
[…]
Wavelength in $x$-direction $\lambda_{x}$ [km] 11
Wavelength in $y$-direction $\lambda_{y}$ [km] $300$ $30$
Wavelength in $z$-direction $\lambda_{z}$ [km] 1-1 3-3 1-1
Branch of intrinsic frequency $\hat{\omega}$ +1+1 +1+1 +1+1
Wave amplitude factor $a_{0}$ $0.5$ $0.5$ $0.1$
Width of wave packet in $x$-direction $\sigma_{x}$ [km] $1500$ $150$
Width of wave packet in $z$-direction $\sigma_{z}$ [km] 55 55 22
Wave packet position in $x$-direction $x_{0}$ [km] $4500$ $400$
Wave packet position in $z$-direction $z_{0}$ [km] $30$ $10$ $10$
Jet strength $u_{0}$ [ms-1] $10$
Width of jet in $z$-direction $z_{0}$ [km] $10$
Jet position in $z$-direction $z_{0}$ [km] $10$
Coriolis parameter $f$ [s-1] $10^{-4}$ $10^{-4}$ 0
Brunt-Vaisala frequency $N$ [s-1] $0.02$ $0.02$ $0.02$
Atmosphere temperature $T_{0}$ [K] 300 300 300
Tracer mixing ratio factor $\alpha_{\psi}$ [m-1] 1 1 1
Domain size in $x$-direction $L_{x}$ [km] $9000$ $800$ 11
Domain size in $y$-direction $L_{y}$ [km] $300$ $30$
Domain size in $z$-direction $L_{z}$ [km] $100$ $100$ $30$
Number of grid points in $x$-direction (WR) $N_{x}$ $512$ $460$ 32
Number of grid points in $y$-direction (WR) $N_{y}$ $16$ $16$
Number of grid points in $z$-direction (WR) $N_{z}$ $1000$ $1000$ 960
Number of grid points in $x$-direction (CR) $N_{x}$ $32$ $32$ 3
Number of grid points in $y$-direction (CR) $N_{y}$ 11 11
Number of grid points in $z$-direction (CR) $N_{z}$ $100$ $100$ 300
In the coarse-resolution simulations 3DCR and 3DCRno, the large-scale tracer distribution $\left\langle{\psi}\right\rangle$ and velocity $\left\langle{\mathbf{v}}\right\rangleare$ obtained directly as output. However, in the high-resolution wave-resolving simulation 3DWR, the output consists of the full variables $\psi$ and $\mathbf{v}$, which include both the large-scale and gravity wave components. Therefore, post-processing is required to extract $\left\langle{\psi}\right\rangle$ and $\left\langle{\mathbf{v}}\right\rangle$. Since the initial wave packet setup in Equation˜86 assumes no dependence on the $y$-coordinate for the large-scale variables, we can extract them by averaging the full variables over one wavelength in the $y$-direction. The gravity wave components are then determined as
@@
The results after 1500 min are visualized in Figure˜1, showing the gravity wave tracer flux convergence $\mathcal{Q}^{(0)}$ and the resulting change in the tracer distribution for 3DWR, 3DCR, and 3DCRno. In the reference simulation 3DWR, the gravity wave packet leads to a decrease in the tracer mixing ratio to the left and an increase to the right of the wave packet amplitude maximum. This behaviour can be explained by the initial setup of $\mathbf{k}_{0}=(0,l_{0},m_{0})$ and $\nabla\left\langle{\psi}\right\rangle=\alpha_{\psi}\mathbf{e}_{z}$, as the wave packet has changed only slightly during the 1500-minute simulation due to its relatively small vertical group velocity. Comparing 3DCR to the reference simulation 3DWR, we find that the fluxes and the resulting changes in the large-scale tracer distribution are well reproduced. To assess the effect of setting $\mathcal{Q}^{(0)}=0$, we compare 3DCR to 3DCRno, where the latter neglects $\mathcal{Q}^{(0)}$. In 3DCRno, the large-scale tracer $\left\langle{\psi}\right\rangle$ is only advected by the wave-induced large-scale wind (not shown). The resulting $\Delta\left\langle{\psi}\right\rangle$ differs significantly from that in 3DCR, and consequently from 3DWR. Thus, we conclude that the changes in $\Delta\left\langle{\psi}\right\rangle$ observed in 3DWR and 3DCR are significantly impacted by $\mathcal{Q}^{(0)}$ and not merely a result of advection due to the large-scale wind $\left\langle{\mathbf{v}}\right\rangle$.
4.2.2 Wave packet in an atmosphere with a jet
The second test case for validating of the computation of $\mathcal{Q}^{(0)}$ initializes a rotating, isothermal ($T_{0}=$300\text{\,}\mathrm{K}$\$) atmosphere with a jet defined by
$\left\langle{\mathbf{v}}\right\rangle(\mathbf{x},t=0)=\begin{cases}\frac{u_{0}}{2}\left\{1+\cos\left[\frac{\pi\left(z-z_{\text{jet}}\right)}{\sigma_{\text{jet}}}\right]\right\}\mathbf{e}_{x}&\text{for}\;|z-z_{\text{jet}}|\leq\sigma_{\text{jet}}\\ 0&\text{else}\;.\end{cases}$ (91)
Because of the Coriolis effect, the jet’s axis rotates counter-clockwise in the $x$-$y$-plane over an $18\text{\,}\mathrm{h}$ period. The large-sacle tracer distribution is initialized following Equation˜87. A gravity-wave packet, specified by Equations˜85 and 86, is introduced with its maximum located at the vertical center of the jet. The simulation parameters are summarized in Table˜2.
Figure˜2 shows the tracer flux convergence $\mathcal{Q}^{(0)}$ and the resulting six-hour change in the large-scale tracer field, $\Delta\left\langle{\psi}\right\rangle$, for the three simulations, 3DWR, 3DCR, and 3DCRno. Computing $\mathcal{Q}^{(0)}$ with MS-GWaM enables the coarse-resolution run (3DCR) to reporduce the $\Delta\left\langle{\psi}\right\rangle$ seen in the wave-resolving reference (3DWR), with only minor differences attributable to resolution. In contrast, when the Stokes drift contribution is omitted (3DCRno), the coarse-resolution simulation exhibits strong deviations from the reference. Even in the presence of background shear flows, including the leading-order gravity-wave impact is essential for accurate tracer-transport simulations.
Taken together with the first test case (Sections˜4.2.1 and 4.2.2), these results demonstrate that the computation of $\mathcal{Q}^{(0)}$ performs reliably across different background flows and initial wavelength settings.